# Fast, accurate and lightweight sequential simulation-based inference using Gaussian locally linear mappings

### v2

@@ 3 Amortized surrogates for likelihoods and posteriors via GLLiM
3:     if $r=0$ or $r=1$ sample $\{\theta_{n}\}_{n=1}^{N}$ iid $\bm{\theta}_{n}\sim\tilde{p}_{r}(\bm{\theta})$ (without MCMC), $n=1,...,N$.
4:     Otherwise for $r\geq 2$ sample $\{\bm{\theta}_{n}\}_{n=1}^{N}$ via $\bm{\theta}_{n}\sim\tilde{p}_{r}(\bm{\theta})$ using Algorithm 2.
5:     Conditionally on every $\bm{\theta}_{n}$, obtain $\{\bm{y}_{n}\}_{n=1}^{N}$ by sampling (implicitly) $\bm{y}_{n}\sim p(\bm{y}\,|\,\bm{\theta}_{n})$ ($n=1,...,N$).
6:     if r = 0 then
7:         Collect $\mathcal{D}_{0}=\{\bm{\theta}_{n},\bm{y}_{n}\}_{n=1}^{N}$.
8:         Train $q_{\tilde{\bm{\phi}}_{0}}(\bm{y}\,|\,\bm{\theta})$ on $\mathcal{D}_{0}$. Update $K$.
9:         Obtain $q_{\bm{\phi}_{0}}(\bm{\theta}\,|\,\bm{y})$.
10:         Set $\tilde{p}_{1}(\bm{\theta})\leftarrow q_{\bm{\phi}_{0}}(\bm{\theta}\,|\,\bm{y}=\bm{y}_{o})$.
11:     else
12:         Collect $\mathcal{D}_{r}=\mathcal{D}_{r-1}\cup\{\bm{\theta}_{n},\bm{y}_{n}\}_{n=1}^{N}\setminus\mathcal{D}_{0}$.
13:         Train $q_{\tilde{\bm{\phi}}_{r}}(\bm{y}\,|\,\bm{\theta})$ on $\mathcal{D}_{r}$. Update $K$.
14:         Obtain $q_{\bm{\phi}_{r}}(\bm{\theta}\,|\,\bm{y})$.
15:         Set $\tilde{p}_{r+1}(\bm{\theta})\leftarrow q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$  ((optional) $\tilde{p}_{r+1}(\bm{\theta})\leftarrow\frac{p(\bm{\theta})}{\tilde{p}_{r}(\bm{\theta})}q_{\bm{\phi}_{r}}(\bm{\theta}\,|\,\bm{y}_{o})$)
16:     end if
17:end for
@@ 4 SeMPLE: sequential learning of likelihoods and posteriors
We have shown how GLLiM can provide amortized estimations of both the likelihood and the posterior distribution at no additional cost, given training data. This GLLiM feature is exploited in our proposed Sequential Mixture Posterior and Likelihood Estimation (SeMPLE), to obtain a sequence of increasingly improved samplers, to more accurately and more efficiently sample from the targeted posterior of observed data $\bm{y}_{o}$. Similarly to SNL (Papamakarios et al., 2019), given a likelihood approximation $q_{\tilde{\bm{\phi}}}(\bm{y}_{o}\,|\,\bm{\theta})$, SeMPLE targets an approximate posterior $\tilde{p}(\bm{\theta})\propto q_{\tilde{\bm{\phi}}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$ with posterior sampling being carried out using a Markov Chain Monte Carlo (MCMC) procedure. Typically, the MCMC scheme used in SNL is slice sampling (Neal, 2003), which has the advantage to be tuning-free but is likely to have difficulties in exploring high dimensional posteriors with complex multimodalities. As a matter of fact, Papamakarios et al. (2019) recommend SNL “at most for tens of parameters that do not have pathologically strong correlations”. In SeMPLE, since GLLiM is able to learn a closed-form approximation of the posterior distribution (in addition to a likelihood function), such posterior is used as a proposal distribution within a Metropolis-Hastings (MH) algorithm. SeMPLE, by using an efficient EM procedure to fit a flexible proposal distribution, is expected to benefit from a MH step that better guides simulations towards more likely parameter values. To ease the comparison with SNL, the latter is recalled in Appendix D, while SeMPLE is summarized in Algorithms 1-2.
@@
Algorithm 1 shows that the targeted approximate posterior for $\bm{\theta}$ is a function denoted with $\hat{p}_{r}(\bm{\theta})\propto q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$ which is sequentially obtained, where $r=0,...,R$ denotes the SeMPLE rounds for the current approximation [^2] We use the term round instead of the most common iteration, as the latter could be confused with the MCMC iterations that are run within a given SeMPLE round.. In the following we denote by $\bm{\theta}^{(r)}$ and $\bm{y}^{(r)}$ a parameter and an observation simulated in round $r$, respectively. At first ($r=0$) $\hat{p}_{0}(\bm{\theta})$ is initialised at the prior $p(\bm{\theta})$. An initial batch of $N$ pairs $\mathcal{D}_{0}=\{\bm{\theta}_{n}^{(0)},\bm{y}_{n}^{(0)}\}_{n=1}^{N}$ produced from the prior-predictive distribution (step 3) constitutes the starting training data for GLLiM, and the parameters $\tilde{\phi}_{0}$ and ${\phi}_{0}$ are obtained via EM within GLLiM to produce the initial surrogate likelihood and posterior $q_{\tilde{\bm{\phi}}_{0}}(\bm{y}\,|\,\bm{\theta})$ and $q_{\bm{\phi}_{0}}(\bm{\theta}\,|\,\bm{y})$ respectively (steps 8-9). The initial surrogate posterior becomes the current target (step 10) $\tilde{p}_{1}(\bm{\theta})\equiv q_{\bm{\phi}_{0}}(\bm{\theta}\,|\,\bm{y}_{o})$. In the next round ($r=1$), the latter can be sampled from without MH (step 3) since $\tilde{p}_{1}(\bm{\theta})$ is defined as a Gaussian mixture, from which it is trivial to sample $N$ parameters $\{\bm{\theta}_{n}^{(1)}\}_{n=1}^{N}$. These $N$ draws are then plugged into the model simulator to produce $\bm{y}_{n}^{(1)}\sim p(\bm{y}\,|\,\bm{\theta}_{n}^{(1)})$ and define $\mathcal{D}_{1}=\{\bm{\theta}_{n}^{(1)},\bm{y}_{n}^{(1)}\}_{n=1}^{N}$ as the new training data (we discard $\mathcal{D}_{0}$ for future rounds as it is deemed too uninformative for training). GLLiM is then fitted to $\mathcal{D}_{1}$ and the corresponding surrogate likelihood and surrogate posterior are obtained. From this step on, the surrogate posterior is employed as a proposal distribution within a tuning-free MH algorithm (Algorithm 2) since, for an arbitrary prior $p(\bm{\theta})$, when $r\geq 2$ we do not explicitly know which distribution is proportional to the product $p(\bm{\theta})q_{\tilde{\phi}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})$ (for $r=1$ we know by construction that the targeted posterior is a Gaussian mixture), except for the special case of Gaussian and Uniform priors (see Papamakarios & Murray, 2016). In Algorithm 2, the acceptance probability of a proposed $\bm{\theta}^{*}$ is (for $r>1$)
@@
Correlated samples from the posterior $\tilde{p}_{r}$ are obtained by first proposing parameters $\bm{\theta}^{*}$ from a Gaussian mixture model. From the current proposal function $q_{\phi_{r-1}}(\bm{\theta}\,|\,\bm{y}_{o})$ which is the mixture
$\sum_{k=1}^{K}\eta_{k}^{(r-1)}(\bm{y}_{o})\mathcal{N}_{l}(\bm{\theta};\bm{A}_{k}^{(r)}\bm{y}_{o}~{}\!\!+\!\!~{}\bm{b}_{k}^{(r-1)},\bm{\Sigma}_{k}^{(r-1)}),$
@@
a component $k^{*}$ is randomly sampled from the set $\{1,...,K\}$ with associated probabilities $\{\eta_{1}^{(r-1)}(\bm{y}_{0}),...,\eta_{K}^{(r-1)}(\bm{y}_{0})\}$. Then $\bm{\theta}^{*}$ is sampled from the selected component $\mathcal{N}_{l}(\bm{\theta};\bm{A}_{k^{*}}^{(r)}\bm{y}+\bm{b}_{k^{*}}^{(r)},\gamma\bm{\Sigma}_{k^{*}}^{(r)})$ by (optionally) slightly inflating its covariance matrix with a factor $\gamma\geq 1$, which we set to $\gamma=1$ (no inflation) or $\gamma\in[1.1,1.2]$ in our experiments. The inflation factor is to ensure that the proposal is sampled from a broad-enough distribution covering the support of the target (no inflation factor is used when $r=1$). Also, note that the $N$ retained draws via MH are $N$ post-burnin draws, where the burnin consists of 100 iterations, and the last accepted draw from round $r$ of SeMPLE is used to initialize the chain at round $r+1$.
@@
Importantly, our MCMC procedure is tuning-free. We use an independence sampler (Robert & Casella, 2004), relying on GLLiM to provide an informative proposal function. This proposal function is trained on draws accepted in previous rounds which are, reasonably, more spread than the currently targeted posterior. In addition, being a mixture model, it is suitable to deal with multimodalities in the target. This greatly simplifies the construction of SeMPLE as no complex tuning is necessary for the MCMC sampler. While $K$ could in principle be set arbitrarily large, in practice this may slow the EM algorithm convergence. In SeMPLE, the value of $K$ is allowed to decrease between rounds in the update $K$ step lines 8 and 13 in Algorithm 1. Mixture components having probabilities $\tilde{\eta}_{k}$ smaller than a threshold are discarded. In most cases we find that setting a starting value of $K=20$ or 30 produces good results, with a threshold for $\tilde{\eta}_{k}$ typically set to 0.005 (multiple hyperboloid, Ornstein-Uhlenbeck, Lotka-Volterra models), 0.03 (Bernoulli GLM model) or simply 0 for no deletion (Two Moons model). This is to prevent having too many irrelevant components, that could be problematic when fitting the mixture to training data produced via MCMC, as the latter may not manage to visit tiny modes. It is also possible to choose the starting value of $K$ in a principled way. In Deleforge et al. (2014), the Bayesian Information Criterion (BIC) is used to select $K$, as we do in our examples, see Appendix B. An appropriate value for $K$ can be learned by repeatedly computing the BIC on realizations of the prior-predictive distribution, before SeMPLE is started. This is an amortized procedure, as the selected values can be reused regardless the specific observed data $\bm{y}_{o}$. The impact on the overall computational budget is negligible. We have also experimented with a version of SeMPLE that learns a corrected posterior (see the optional part in line 16 of Algorithm 1) using the correction formula from Papamakarios & Murray (2016). However, we found that learning a surrogate likelihood was much more beneficial, providing better inference coupled with a higher acceptance rate. Notice that none of our reported results use the optional corrected-posterior approach.
@@ 5 Numerical illustrations
We illustrate SeMPLE on several simulation studies, see also the Appendices for extra results and additional models. SeMPLE is written in R and uses the GLLiM implementation from the xLLiM package (Perthame et al., 2022). Comparisons with neural-based likelihood and posterior estimation methods (SNL and SNPE-C) use the Python implementations in the SBIBM package for SBI benchmarking (Lueckmann et al., 2021), and architecture specifications for SBIBM are in Appendix A. To compare approximate posteriors with a suitable reference posterior, we consider the classifier two-samples test (C2ST), as implemented in SBIBM, and Wasserstein distances via the POT Python package (Flamary et al., 2021). C2ST varies in the interval [0.5,1], the lower the better. It equals 0.5 when samples from two distributions appear statistically indistinguishable and equals 1 when the classifier is able to perfectly separate samples as belonging to two different distributions. We also report runtimes in figures and resource requirements in Table 1. For SeMPLE, the time required to select $K$ via BIC is reported separately and not taken into account in the figures below. For some case studies, such as Lotka-Volterra, we merely picked a starting value of $K$ without running the BIC procedure. Adding this overhead time does not change the overall conclusion that SeMPLE is much more time-efficient than the other methods. Moreover, the selection of $K$ is an amortized procedure, hence it does not need to be repeated for different observations of the same model.
@@
In the next two sections 5.1 and 5.2, we illustrate results from two models having multimodal posteriors, in section 5.3 we summarise the Lotka-Volterra experiment, and in section 5.5 and Table 2 we summarise results from all other experiments detailed in the Appendices. For comparison, all methods are given the same budget of model simulations. For SNL and SNPE-C the SBIBM default of distributing these simulations uniformly across $R=10$ rounds is used, but to ease comparison we also run SNL and SNPE-C with the same number of rounds used for SeMPLE, which is $R=4$, except for the Bernoulli GLM model where $R=2$ suffices. In fact, we observe that SeMPLE does not need as many sequential steps as the other methods. It may be argued that SNL could perform better if a larger number of MCMC iterations was used, at the cost of a smaller number of rounds $R$. Results in the Appendices show that both SNL and SNPE-C perform worse under this alternative configuration. Regarding SeMPLE, when the computational budget limits the number of model simulations, we found it beneficial to keep the number of rounds $R$ between 2 and 3, where $R=2$ implies that we have exactly one round where MCMC is used, and this is typically already enough for satisfying inference, though usually an additional round with MCMC (hence $R=3$) produces further refined inference. In our work we show results up to $R=4$, but in general, four rounds are not recommended. The algorithm performance relies first on the GLLiM ability to learn reasonably accurate likelihood and posterior approximations and then on the MCMC sampler to explore the parameter space. Both steps require a large enough number of model simulations, which depends mainly on the model dimension.

### v3

@@ 3 Amortized surrogates for likelihoods and posteriors via GLLiM
3:     if $r=0$ or $r=1$ then
4:         sample $\{\bm{\theta}_{n}\}_{n=1}^{N}$ iid $\bm{\theta}_{n}\sim\tilde{p}_{r}(\bm{\theta})$ (without MCMC), $n=1,...,N$.
5:     else
6:         sample $\{\bm{\theta}_{n}\}_{n=1}^{N}$ via $\bm{\theta}_{n}\sim\tilde{p}_{r}(\bm{\theta})$ using Algorithm 2.
7:     end if
8:     Conditionally on every $\bm{\theta}_{n}$, obtain $\{\bm{y}_{n}\}_{n=1}^{N}$ by sampling (implicitly) $\bm{y}_{n}\sim p(\bm{y}\,|\,\bm{\theta}_{n})$ ($n=1,...,N$).
9:     if r = 0 then
10:         Collect $\mathcal{D}_{0}=\{\bm{\theta}_{n},\bm{y}_{n}\}_{n=1}^{N}$.
11:         Train $q_{\tilde{\bm{\phi}}_{0}}(\bm{y}\,|\,\bm{\theta})$ on $\mathcal{D}_{0}$. Update $K$.
12:         Obtain $q_{\bm{\phi}_{0}}(\bm{\theta}\,|\,\bm{y})$.
13:         Set $\tilde{p}_{1}(\bm{\theta})\leftarrow q_{\bm{\phi}_{0}}(\bm{\theta}\,|\,\bm{y}=\bm{y}_{o})$.
14:     else
15:         Collect $\mathcal{D}_{r}=\mathcal{D}_{r-1}\cup\{\bm{\theta}_{n},\bm{y}_{n}\}_{n=1}^{N}\setminus\mathcal{D}_{0}$.
16:         Train $q_{\tilde{\bm{\phi}}_{r}}(\bm{y}\,|\,\bm{\theta})$ on $\mathcal{D}_{r}$. Update $K$.
17:         Obtain $q_{\bm{\phi}_{r}}(\bm{\theta}\,|\,\bm{y})$.
18:         Set $\tilde{p}_{r+1}(\bm{\theta})\leftarrow q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$
19:          ((optional) $\tilde{p}_{r+1}(\bm{\theta})\leftarrow\frac{p(\bm{\theta})}{\tilde{p}_{r}(\bm{\theta})}q_{\bm{\phi}_{r}}(\bm{\theta}\,|\,\bm{y}_{o})$)
20:     end if
21:end for
@@ 4 SeMPLE: sequential learning of likelihoods and posteriors
We have shown how GLLiM can provide amortized estimations of both the likelihood and the posterior distribution at no additional cost, given training data. This GLLiM feature is exploited in our proposed Sequential Mixture Posterior and Likelihood Estimation (SeMPLE), to obtain a sequence of increasingly improved samplers, to more accurately and more efficiently sample from the targeted posterior of observed data $\bm{y}_{o}$. Similarly to SNL (Papamakarios et al., 2019), given a likelihood approximation $q_{\tilde{\bm{\phi}}}(\bm{y}_{o}\,|\,\bm{\theta})$, SeMPLE targets an approximate posterior $\tilde{p}(\bm{\theta})\propto q_{\tilde{\bm{\phi}}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$ with posterior sampling being carried out using a MCMC procedure. Typically, the MCMC scheme used in SNL is slice sampling (Neal, 2003), which has the advantage to be tuning-free but is likely to have difficulties in exploring high dimensional posteriors with complex multimodalities. As a matter of fact, Papamakarios et al. (2019) recommend SNL “at most for tens of parameters that do not have pathologically strong correlations”. In SeMPLE, since GLLiM is able to learn a closed-form approximation of the posterior distribution (in addition to a likelihood function), such posterior is used as a proposal distribution within a Metropolis-Hastings (MH) algorithm. SeMPLE, by using an efficient EM procedure to fit a flexible proposal distribution, is expected to benefit from a MH step that better guides simulations towards more likely parameter values. To ease the comparison with SNL, the latter is recalled in Appendix D, while SeMPLE is summarized in Algorithms 1-2.
@@
Algorithm 1 shows that the targeted approximate posterior for $\bm{\theta}$ is a function denoted with $\hat{p}_{r}(\bm{\theta})\propto q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$ which is sequentially obtained, where $r=0,...,R$ denotes the SeMPLE rounds for the current approximation [^3] We use the term round instead of the most common iteration, as the latter could be confused with the MCMC iterations that are run within a given SeMPLE round.. In the following we denote by $\bm{\theta}^{(r)}$ and $\bm{y}^{(r)}$ a parameter and an observation simulated in round $r$, respectively. At first ($r=0$) $\hat{p}_{0}(\bm{\theta})$ is initialised at the prior $p(\bm{\theta})$. An initial batch of $N$ pairs $\mathcal{D}_{0}=\{\bm{\theta}_{n}^{(0)},\bm{y}_{n}^{(0)}\}_{n=1}^{N}$ produced from the prior-predictive distribution (step 4) constitutes the starting training data for GLLiM, and the parameters $\tilde{\phi}_{0}$ and ${\phi}_{0}$ are obtained via EM within GLLiM to produce the initial surrogate likelihood and posterior $q_{\tilde{\bm{\phi}}_{0}}(\bm{y}\,|\,\bm{\theta})$ and $q_{\bm{\phi}_{0}}(\bm{\theta}\,|\,\bm{y})$ respectively (steps 11-12). The initial surrogate posterior becomes the current target (step 13) $\tilde{p}_{1}(\bm{\theta})\equiv q_{\bm{\phi}_{0}}(\bm{\theta}\,|\,\bm{y}_{o})$. In the next round ($r=1$), the latter can be sampled from without MH (step 4) since $\tilde{p}_{1}(\bm{\theta})$ is defined as a Gaussian mixture, from which it is trivial to sample $N$ parameters $\{\bm{\theta}_{n}^{(1)}\}_{n=1}^{N}$. These $N$ draws are then plugged into the model simulator to produce $\bm{y}_{n}^{(1)}\sim p(\bm{y}\,|\,\bm{\theta}_{n}^{(1)})$ and define $\mathcal{D}_{1}=\{\bm{\theta}_{n}^{(1)},\bm{y}_{n}^{(1)}\}_{n=1}^{N}$ as the new training data (we discard $\mathcal{D}_{0}$ for future rounds as it is deemed too uninformative for training). GLLiM is then fitted to $\mathcal{D}_{1}$ and the corresponding surrogate likelihood and surrogate posterior are obtained. From this step onwards, the surrogate posterior is employed as a proposal distribution within a tuning-free MH algorithm (Algorithm 2) since, for an arbitrary prior $p(\bm{\theta})$, when $r\geq 2$ we do not explicitly know which distribution is proportional to the product $p(\bm{\theta})q_{\tilde{\phi}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})$ (for $r=1$ we know by construction that the targeted posterior is a Gaussian mixture), except for the special case of Gaussian and Uniform priors (see Papamakarios & Murray, 2016). In Algorithm 2, the acceptance probability of a proposed $\bm{\theta}^{*}$ is (for $r>1$)
@@
Samples from the posterior $\tilde{p}_{r}$ are obtained by first proposing parameters $\bm{\theta}^{*}$ from a Gaussian mixture model. From the current proposal function $q_{\phi_{r-1}}(\bm{\theta}\,|\,\bm{y}_{o})$, which is the mixture
$q_{\phi_{r-1}}(\bm{\theta}\,|\,\bm{y}_{o})=\sum_{k=1}^{K}\eta_{k}^{(r-1)}(\bm{y}_{o})\mathcal{N}_{l}(\bm{\theta};\bm{A}_{k}^{(r-1)}\bm{y}_{o}~{}\!\!+\!\!~{}\bm{b}_{k}^{(r-1)},\bm{\Sigma}_{k}^{(r-1)}),$
@@
a component $k^{*}$ is randomly sampled from the set $\{1,...,K\}$ with associated probabilities $\{\eta_{1}^{(r-1)}(\bm{y}_{0}),...,\eta_{K}^{(r-1)}(\bm{y}_{0})\}$. Then $\bm{\theta}^{*}$ is sampled from the $k^{*}$-th component $\mathcal{N}_{l}(\bm{\theta};\bm{A}_{k^{*}}^{(r-1)}\bm{y}+\bm{b}_{k^{*}}^{(r-1)},\gamma\bm{\Sigma}_{k^{*}}^{(r-1)})$ by (optionally) slightly inflating its covariance matrix with a factor $\gamma\geq 1$, which we set to $\gamma=1$ (no inflation) or $\gamma\in[1.1,1.2]$ in our experiments. The inflation factor is to ensure that the proposal is sampled from a broad-enough distribution covering the support of the target (no inflation factor is used when $r=1$). Also, note that the $N$ retained draws via MH are $N$ post-burnin draws, where the burnin consists of 100 iterations, and the last accepted draw from round $r$ of SeMPLE is used to initialize the chain at round $r+1$.
@@
Importantly, our MCMC procedure is tuning-free. We use an independence sampler (Robert & Casella, 2004), relying on GLLiM to provide an informative proposal function. This proposal function is trained on draws accepted in previous rounds which are, reasonably, more spread than the currently targeted posterior. In addition, being a mixture model, it is suitable to deal with multimodalities in the target. This greatly simplifies the construction of SeMPLE as no complex tuning is necessary for the MCMC sampler. While $K$ could in principle be set arbitrarily large, in practice this may slow the EM algorithm convergence. In SeMPLE, the value of $K$ is allowed to decrease between rounds, as expressed with “update $K$ in lines 11 and 16 of Algorithm 1. In fact, mixture components having probabilities $\tilde{\eta}_{k}$ smaller than a threshold are discarded (ie these gets removed from the mixture model). In most cases we find that setting a starting value of $K=20$ or 30 produces good results, with a threshold for $\tilde{\eta}_{k}$ typically set to 0.005 (multiple hyperboloid, Ornstein-Uhlenbeck, Lotka-Volterra, biological SDE model), 0.03 (Bernoulli GLM model) or simply 0 for no deletion (Two Moons model). This is to prevent having too many irrelevant components, that could be problematic when fitting the mixture to training data produced via MCMC, as the latter may not manage to visit tiny modes. It is also possible to choose the starting value of $K$ in a principled way. In Deleforge et al. (2014), the Bayesian Information Criterion (BIC) is used to select $K$, as we do in some of our examples, see Appendix B. An appropriate value for $K$ can be learned by repeatedly computing the BIC on realizations of the prior-predictive distribution, before SeMPLE is started. This use of the BIC is an amortized procedure, as the selected values can be reused regardless the specific observed data $\bm{y}_{o}$. The impact on the overall computational budget is negligible. We have also experimented with a version of SeMPLE that learns a corrected posterior (see the optional part in step 19 of Algorithm 1) using the correction formula from Papamakarios & Murray (2016). However, we found that learning a surrogate likelihood was much more beneficial, providing better inference coupled with a higher acceptance rate. Notice that none of our reported results use the optional corrected-posterior approach.
@@ 5 Numerical illustrations
We illustrate SeMPLE on several simulation studies, see also the Appendices for extra results and additional models. SeMPLE is written in R and uses the GLLiM implementation from the xLLiM package (Perthame et al., 2022). Comparisons with neural-based likelihood and posterior estimation methods (SNL and SNPE-C) use the Python implementations in the SBIBM package for SBI benchmarking (Lueckmann et al., 2021), and architecture specifications for SBIBM are in Appendix A. To compare approximate posteriors with a suitable reference posterior, we consider the classifier two-samples test (C2ST), as implemented in SBIBM, and Wasserstein distances via the POT Python package (Flamary et al., 2021). C2ST varies in the interval [0.5,1], the lower the better. It equals 0.5 when samples from two distributions appear statistically indistinguishable and equals 1 when the classifier is able to perfectly separate samples as belonging to two different distributions. We also report runtimes in figures and resource requirements in Table 1. For SeMPLE, the time required to select $K$ via BIC (when the BIC procedure is performed, which is not the case for all studies) is reported separately and not taken into account in the figures below. In fact, for some case studies, such as Lotka-Volterra, we merely picked a starting value of $K$ without running the BIC procedure. When we used the BIC, adding this overhead time does not change the overall conclusion that SeMPLE is much more time-efficient than the other methods. Moreover, the selection of $K$ is an amortized procedure, hence it does not need to be repeated for different observations of the same model.
@@
In the next two sections 5.1 and 5.2, we illustrate results from two models having multimodal posteriors, in section 5.3 we summarise the Lotka-Volterra experiment, and in section 5.6 and Table 2 we summarise results from other experiments detailed in the Appendices. For comparison, all methods are given the same budget of model simulations. For SNL and SNPE-C the SBIBM default of distributing these simulations uniformly across $R=10$ rounds is used, but to ease comparison we also run SNL and SNPE-C with the same number of rounds used for SeMPLE, which is $R=4$, except for the Bernoulli GLM model where $R=2$ suffices. In fact, we observe that SeMPLE does not need as many sequential steps as the other methods. It may be argued that SNL could perform better if a larger number of MCMC iterations was used, at the cost of a smaller number of rounds $R$. Results in the Appendices show that both SNL and SNPE-C perform worse under this alternative configuration. Regarding SeMPLE, when the computational budget limits the number of model simulations, we found it beneficial to keep the number of rounds $R$ between 2 and 3, where $R=2$ implies that we have exactly one round where MCMC is used, and this is typically already enough for satisfying inference, though usually an additional round with MCMC (hence $R=3$) produces further refined inference. In our work we show results up to $R=4$, but in general, four rounds are not recommended. The algorithm performance relies first on the GLLiM ability to learn reasonably accurate likelihood and posterior approximations and then on the MCMC sampler to explore the parameter space. Both steps require a large enough number of model simulations, which depends mainly on the model dimension.